# BEVPlace++: Fast, Robust, and Lightweight LiDAR Global Localization for Unmanned Ground Vehicles

### v2

@@ Abstract
This article introduces BEVPlace++, a novel, fast, and robust LiDAR global localization method for unmanned ground vehicles. It uses lightweight convolutional neural networks (CNNs) on Bird’s Eye View (BEV) image-like representations of LiDAR data to achieve accurate global localization through place recognition followed by 3-DoF pose estimation. Our detailed analyses reveal an interesting fact that CNNs are inherently effective at extracting distinctive features from LiDAR BEV images. Remarkably, keypoints of two BEV images with large translations can be effectively matched using CNN-extracted features. Building on this insight, we design a rotation equivariant module (REM) to obtain distinctive features while enhancing robustness to rotational changes. A Rotation Equivariant and Invariant Network (REIN) is then developed by cascading REM and a descriptor generator, NetVLAD, to sequentially generate rotation equivariant local features and rotation invariant global descriptors. The global descriptors are used first to achieve robust place recognition, and the local features are used for accurate pose estimation. Experimental results on multiple public datasets demonstrate that BEVPlace++, even when trained on a small dataset (3000 frames of KITTI) only with place labels, generalizes well to unseen environments, performs consistently across different days and years, and adapts to various types of LiDAR scanners. BEVPlace++ achieves state-of-the-art performance in subtasks of global localization including place recognition, loop closure detection, and global localization. Additionally, BEVPlace++ is lightweight, runs in real-time, and does not require accurate pose supervision, making it highly convenient for deployment. The source codes are publicly available at https://github.com/zjuluolun/BEVPlace.
@@ I Introduction
Global localization aims to estimate the global poses of robots on a map without other prior information, which is a key component for achieving full autonomy of unmanned ground vehicles and crucial for many robotic applications. Especially in Simultaneous Localization and Mapping (SLAM) [1, 2, 3, 4], global localization provides loop closure constraints to eliminate accumulative drifts, essential for building globally consistent maps. Furthermore, it generates complementary pose estimations, which is important for initializing pose tracking or recovering it from failures [5, 6].
@@
A widely adopted global localization paradigm typically first builds a global map with structure from motion or SLAM. When a query image or LiDAR scan is received, the system searches the map to find the best-matched place and then computes the global pose through sensor data registration. Over the past two decades, image-based global localization methods [7, 8, 2, 9, 10, 11] have been well-developed, thanks to the strong image features such as ORB [12] and SIFT [13]. Despite their broad usage in robotics, image-based methods are not robust enough to illumination changes and are sensitive to view changes due to the perspective imaging mechanism. In contrast, LiDARs actively emit infrared beams and measure the time of flight to perceive depth information, making LiDAR-based localization naturally robust to lighting changes [14]. Additionally, precise depth information provided by LiDARs allows for more accurate localization results. In the early days, LiDAR-based frameworks were less favored due to the immature manufacturing process of LiDARs. However, advancements in technology have made modern LiDARs more accurate, smaller, and cheaper, leading to their widespread application in fields such as mobile robots and autonomous driving. Thus, LiDAR-based global localization has become a hot topic and a crucial problem for robot perception tasks.
Despite the great potential of LiDAR in global localization, the sparsity of point clouds presents a significant challenge. Unlike image pixels organized in regular grids, point clouds are sparsely distributed in continuous 3D space, making it difficult to extract stable keypoints and distinctive features, which can lead to poor localization accuracy. Additionally, point clouds are typically dense for nearby scenes but sparse for distant ones, with varying sparsity across different types of LiDAR scanners. These factors complicate the generalization of LiDAR-based localization methods to different sensor configurations and new environments.
Some previous works tackle LiDAR localization by projecting point clouds into range images [15, 16, 17, 18]. Using range images avoids processing the sparse point clouds and shows promising performance across different environments. Moreover, the rotation of point clouds is equivalent to the horizontal shifting of range images, making range image-based methods robust to view rotations. However, range images suffer from scale distortions due to the spherical projection, limiting localization accuracy. Other methods try to use point networks to learn distinctive point features [19, 20, 21]. With accurate pose supervision, these approaches can achieve high localization accuracy but may not adapt well to different LiDAR scanner configurations or environments. Some methods build local maps from consecutive LiDAR scans and perform localization using these local maps [22, 23, 24]. While local maps can reduce the impact of sparsity, constructing them requires the robot to travel a certain distance and properly align the local map, which may compromise safety and is not always feasible.
@@
This work introduces BEVPlace++, a fast, robust, and lightweight LiDAR global localization method for 3-DoF pose estimation without knowing the initial pose [25]. BEVPlace++ uses Bird’s Eye View (BEV) images as a simple yet effective and lightweight representation of LiDAR data. Unlike range images, BEV images provide more stable object scales and relationships, offering better generalization and easier deployment across different LiDAR sensors. By leveraging the translation equivariance of convolutional neural networks (CNNs) [26], we show that CNN features of BEV images are inherently distinctive and robust to the sparsity of LiDAR point clouds, allowing for accurate pose estimation between BEV image pairs. Such properties enable BEVPlace++ to generalize to unseen environments and adapt to different types of LiDAR scanners in a zero-shot learning manner. Additionally, we design a rotation equivariant Module (REM) to enhance BEVPlace++’s robustness to view changes. Both theoretical and statistical validations confirm the rotation equivariance of our devised REM, enabling BEV images with large view changes to be matched. Building based on REM, we propose a Rotation Equivariant and Invariant Network (REIN) by cascading REM with NetVLAD [27] to generate rotation equivariant local features and rotation invariant global descriptors. Unlike previous works requiring accurate pose labels [19, 28, 20, 21], BEVPlace++ achieves both place recognition and pose estimation using only coarse positive/negative place labels for supervision. An overview of the proposed method is depicted in Fig. 1. We use triplet supervision to minimize the feature distance of similar BEV images and maximize the distance of dissimilar ones. Benefiting from our designed translation and rotation equivariant network, such place supervision is sufficient to extract distinctive features on BEV images, allowing direct accurate pose estimation with Random Sample Consensus (RANSAC) [29] matching between BEV images.
In summary, the contributions of this work are five folds:
  • We propose BEVPlace++, a complete global localization framework that utilizes the simple yet effective BEV representation of point clouds for both place recognition and pose estimation.
  • We provide a detailed analysis of how the translation equivariance of CNN helps BEV image matching. Our BEVPlace++ accurately aligns BEV images with large translational differences using CNN features and effectively adapts to new, unseen environments.
  • We propose a novel Rotation Equivariant and Invariant Network (REIN), constructed with lightweight CNN blocks and NetVLAD. It helps BEVPlace++ achieve robust place recognition and pose estimation under significant rotational view changes.
  • We conducted comprehensive experiments on five different public datasets spanning various days, years, environments, and LiDAR scanner types. Our results demonstrate that BEVPlace++ achieves state-of-the-art performance across multiple tasks, including place recognition, loop closing, global localization, and SLAM.
  • We open-source concise Python APIs of BEVPlace++, contributing to the robotics community.
@@
BEVPlace++ is an extension of our previous conference paper BEVPlace [30]. BEVPlace uses group convolution to extract rotation equivariant local features from BEV images and achieves place recognition by global descriptor matching. Compared to BEVPlace, BEVPlace++ extends in four critical aspects: 1) a novel REIN network, with a faster and more light-weighted rotation equivariant and invariant feature encoder; 2) a deeper analysis of the rotation and translation equivariance of CNN features; 3) a complete pose estimator for achieving 3-DoF poses global localization; 4) more extensive and comprehensive experimental evaluations on various datasets, demonstrating superior performance in place recognition, loop closure detection, and complete global localization.
@@ II Related Work
This section briefly overviews LiDAR-based global localization methods in the literature. Following existing surveys [31, 32], we categorize related works into three groups based on the integration degree of place recognition and pose estimation: 1) place recognition-only, that retrieves the most similar place using global descriptors; 2) global pose estimation, that directly estimates the global poses without retrieving places; and 3) place recognition followed by local pose estimation, that first achieves place recognition and then estimates the robot pose via a pose estimator.
@@
1) Place recognition-only. These methods use a divide-and-conquer strategy, generating global descriptors for LiDAR scans and performing place recognition through descriptor retrieval. The pose of the retrieved place is then considered as the estimated pose. The key to place recognition is creating global descriptors that ensure similar scans are close in descriptor space while dissimilar ones are far apart. Early methods usually exploit the point statistics to represent the point cloud appearance. For example, M2DP [33] projects a point cloud to multiple 2D planes and generates a density signature for points in each plane. The singular value decomposition (SVD) components of the signature are then used to compute a global descriptor. Scan Context [34, 35] partitions the ground space into bins according to both azimuthal and radial directions and encodes the 2.5-D information within an image.
@@
Recently, learning-based place recognition has been a hot topic, and many methods have been proposed. PointNetVLAD [14] leverages a network to project each point into a higher dimension feature and then uses NetVLAD [27] to generate global features. To take advantage of more contextual information, PCAN [36] introduces the point contextual attention network that learns attention to task-relevant features. Both PointNetVLAD and PCAN cannot capture local geometric structures due to the independent treatment for each point. Thus, the following methods focus on extracting more distinctive local features considering the neighborhood information. LPD-Net [37] adopts an adaptive local feature module to extract the handcrafted features and uses a graph-based neighborhood aggregation module to discover the spatial distribution of local features. EPC-Net [38] improves LPD-Net by using a proxy point convolutional neural network. SOE-Net [39] introduces a point orientation encoding (PointOE) module. Minkloc3D [40, 41] uses sparse 3D convolutions in local areas and achieves state-of-the-art performance on the benchmark dataset. Recently, some works including SVT-Net [42], TransLoc3D [43], NDT-Transformer [44], and PPT-Net [45] leverage the transformer-based attention mechanism [46] to boost place recognition performance. However, all these data-driven methods may lack generalization ability to unseen environments. Some methods explore the potential of the range image representation of point clouds. OverlapNet [15] uses the overlap of range images to determine whether two point clouds are at the same place and uses a Siamese network to estimate the overlap. OverlapTransformer [17] further uses a transformer architecture to learn rotation-invariant global features. CVTNet [18] combines range images and BEV images to perform matching. It transforms BEV images into a format similar to range images to achieve rotation-invariance. The range image-based methods usually have better generalization ability but suffer from the scale distortions of translation movements of point clouds.
@@
2) Global pose estimation. These methods utilize the local characteristics of point clouds through geometric measures to generate local descriptors, which are then directly matched with a global map to determine the global poses. For example, the fast point feature histograms (FPFH) [47] uses the relationship between the neighbors of interest points and encodes them into a histogram. The Signature of Histograms of OrienTations (SHOT) [48] builds a local reference frame and leverages the orientation distribution of the normals in local regions. Although these methods can align local LiDAR scans, they usually lack descriptive capability in outdoor environments due to their sensitivity to point cloud density and noise. To leverage the local descriptors more efficiently, the keypoints voting method [49] performs localization using the 3D Gestalt descriptors. This method relies on robust keypoint extraction, while repeatable 3D keypoint detection is still an open problem in the literature. Some studies [50, 51] directly extract conventional image descriptors such as ORB [12] and SIFT [13] from LiDAR images for place recognition. These methods show promising performance in small environments but usually lack enough distinctiveness in large-scale outdoor environments. SegMatch [22] performs segmentation on LiDAR scans and builds a segment map. During online global localization, it extracts segments from the query LiDAR scan and matches them with the map. The following SegMap [23] further assigns the segment descriptors learned by a neural network to improve the matching performance. The main drawback of these methods is their reliance on traveling relatively long distances to extract distinctive segments by gathering multiple scans.
@@
3) Place recognition followed by local pose estimation. These approaches have been commonly used in visual global localization. However, it is not widely adopted in LiDAR global localization because few LiDAR local features have reached the same level of maturity as visual features like ORB and SIFT. To tackle the problem, BVMatch [52] projects point clouds into BEV images and extracts handcrafted BVFT features from the images. It then uses the bag-of-words model [7, 8, 11] to generate global features. However, it is shown that BVMatch cannot generalize well to unseen environments. BoW3D [53] extracts local features directly from sparse point clouds. It also adopts the bag-of-words model to generate global descriptors and is expected to show moderate generalization ability. Another line of work adopts deep learning techniques. For example, DH3D [19] uses the 3D local feature encoder and detector to extract local descriptors. It embeds the descriptors to a global feature for place recognition and aligns the matched LiDAR pairs using RANSAC [29]. The following methods such as LCDNet [20], LoGG3D-Net [28], and LCRNet [21] use a similar strategy and unify place recognition and pose estimation into one framework. LCDNet extracts distinctive local features with pointvoxel-RCNN [54] and performs data association with the Sinkhorn algorithm [55]. LoGG3D-Net [28] introduces a local consistency loss to guide the network toward learning local features consistently across revisits. LCRNet [21] exploits novel feature extraction and pose-aware attention mechanism to precisely estimate similarities and 6-DoF poses between pairs of LiDAR scans. By jointly learning local and global descriptors, these methods show satisfactory global localization performance. However, they may not adapt well when the point clouds are out of the distribution of the training data. BTC [24] extracts the keypoints of a point cloud by projecting the points to planes. It generates triangle descriptors and develops an efficient matching strategy. However, the BTC descriptor suffers from sparsity of point clouds as keypoints detection is unstable in sparse single LiDAR scans.
BEVPlace++ follows a two-step paradigm, sequentially performing place recognition and pose estimation for global localization. By first performing coarse localization through place recognition, unrelated scenes or BEV images can be effectively filtered out. In dynamic environments, the place recognition stage can help adapt to changes in the scene, such as lighting variations or partial occlusions. Unlike previous two-step-based methods, BEVPlace++ uses the BEV representation of point clouds. By leveraging the translation and rotation equivariance of CNNs, BEVPlace++ achieves robust pose estimation without requiring accurate pose supervision. Based on our devised REIN network, BEVPlace++ demonstrates remarkable generalization in unseen environments and adapts effectively to various sensor configurations.

### v3

@@ Abstract
This article introduces BEVPlace++, a novel, fast, and robust LiDAR global localization method for unmanned ground vehicles. It uses lightweight convolutional neural networks (CNNs) on Bird’s Eye View (BEV) image-like representations of LiDAR data to achieve accurate global localization through place recognition, followed by 3-DoF pose estimation. Our detailed analyses reveal an interesting fact that CNNs are inherently effective at extracting distinctive features from LiDAR BEV images. Remarkably, keypoints of two BEV images with large translations can be effectively matched using CNN-extracted features. Building on this insight, we design a Rotation Equivariant Module (REM) to obtain distinctive features while enhancing robustness to rotational changes. A Rotation Equivariant and Invariant Network (REIN) is then developed by cascading REM and a descriptor generator, NetVLAD, to sequentially generate rotation equivariant local features and rotation invariant global descriptors. The global descriptors are used first to achieve robust place recognition, and then local features are used for accurate pose estimation. Experimental results on seven public datasets and our UGV platform demonstrate that BEVPlace++, even when trained on a small dataset (3000 frames of KITTI) only with place labels, generalizes well to unseen environments, performs consistently across different days and years, and adapts to various types of LiDAR scanners. BEVPlace++ achieves state-of-the-art performance in multiple tasks, including place recognition, loop closure detection, and global localization. Additionally, BEVPlace++ is lightweight, runs in real-time, and does not require accurate pose supervision, making it highly convenient for deployment. The source codes are publicly available at https://github.com/zjuluolun/BEVPlace2.
@@ I Introduction
Global localization estimates the global poses of robots on a map without prior information, playing a pivotal role in enabling full autonomy for unmanned ground vehicles and supporting a range of robotic applications. For example, in Simultaneous Localization and Mapping (SLAM) [1, 2, 3, 4], global localization provides loop closure constraints that reduce accumulative drifts and ensure globally consistent maps. It also offers complementary pose estimations crucial for initializing or recovering pose tracking [5, 6].
@@
A widely adopted global localization paradigm first builds a global map using structure-from-motion or SLAM. Upon receiving a query image or LiDAR scan, the method searches the map for the best-matched place and computes the global pose through sensor data registration. While image-based global localization methods [7, 8, 2, 9, 10, 11] have been extensively developed, leveraging robust features such as ORB [12] and SIFT [13], they are sensitive to illumination and view changes due to perspective imaging. In contrast, LiDAR-based methods, benefiting from active depth sensing, are naturally robust to lighting variations [14] and provide precise depth information for accurate localization. The evolution of LiDAR technology has made sensors more accurate, compact, and cost-effective, making LiDAR-based global localization an important topic in 3D visual SLAM and driving its widespread adoption in mobile robotics and autonomous vehicles.
Despite the great potential of LiDAR global localization, it faces significant challenges in generalization ability. The reasons are twofold. First, since LiDAR scanners detect depth under fixed angular resolutions, point clouds are correspondingly dense for nearby scenes but sparse for distant ones, with varying sparsity across different types of LiDAR scanners. This makes it difficult to extract stable keypoints and distinctive features, which can lead to poor localization accuracy. Second, current learning-based global localization methods  [15, 16, 17, 18] usually need pose supervision. As accurate ground truth poses are expensive to obtain, these methods are difficult to train and may hardly generalize to point clouds collected from different types of sensors.
@@
This work introduces BEVPlace++, a fast, robust, and lightweight LiDAR global localization method for 3-DoF global pose estimation without knowing the initial pose. To relieve the influence of the sparsity of point clouds, BEVPlace++ uses Bird’s Eye View (BEV) images as a simple yet effective and lightweight representation. Unlike the broadly used range images [19, 20, 21, 22], BEV images provide more stable object scales and relationships, offering better generalization and easier deployment across different LiDAR sensors. We further show that CNN features of BEV images are inherently distinctive and robust to the sparsity of LiDAR point clouds, allowing for accurate pose estimation between BEV image pairs without accurate pose supervision. Additionally, we design a rotation equivariant Module (REM) to enhance BEVPlace++’s robustness to view changes. Both theoretical and statistical validations confirm the rotation equivariance of our devised REM, enabling BEV images with large view changes to be matched. Building upon REM, we propose a Rotation Equivariant and Invariant Network (REIN) by cascading REM with NetVLAD [23] to generate rotation equivariant local features and rotation invariant global descriptors. Unlike previous works requiring accurate pose labels [15, 24, 17, 18], BEVPlace++ achieves both place recognition and pose estimation using only coarse positive/negative place labels for supervision. An overview of the proposed method is depicted in Fig. 1. We use triplet supervision to minimize the feature distance of similar BEV images and maximize the distance of dissimilar ones. Benefiting from our designed translation and rotation equivariant network, such place supervision is sufficient to extract distinctive features on BEV images, allowing direct accurate pose estimation with Random Sample Consensus (RANSAC) [25] matching between BEV images.
In summary, the contributions of this work are four folds:
  • We introduce a weakly supervised BEV image-based LiDAR global localization framework named BEVPlace++. Exhaustive experiments on 7 public datasets demonstrate its state-of-the-art performance in place recognition, loop closure, global localization, and SLAM, with strong generalization to diverse environments and sensor types.
  • We provide an empirical analysis showing that deep CNN features of BEV images are naturally discriminative for local feature matching under translation motions. Based on this observation, we propose a Rotation Equivariant Module (REM) to preserve the distinctiveness of deep features under rotations by directly leveraging CNN backbones within the module.
  • We propose a novel Rotation Equivariant and Invariant Network (REIN) combining REM and NetVLAD. REIN makes BEVPlace++ robust to yaw angle view changes by extracting rotation-invariant global descriptors for place recognition and local rotation-equivariant descriptors for pose estimation. Additionally, REIN does not require accurate pose supervision, relying instead on coarse place information for training.
  • We heve open-sourced BEVPlace++ contributing to the community at https://github.com/zjuluolun/BEVPlace2.
@@
BEVPlace++ is an extension of our previous conference paper BEVPlace [26]. BEVPlace uses group convolution to extract rotation equivariant local features from BEV images and achieves place recognition by global descriptor matching. Compared to BEVPlace, BEVPlace++ extends in four critical aspects: 1) a novel REIN network, with a faster and more light-weight rotation equivariant and invariant feature encoder; 2) a deeper analysis of the rotation and translation equivariance of CNN features; 3) a pose estimator for achieving 3-DoF poses global localization; 4) more extensive and comprehensive experimental evaluations on various datasets, demonstrating superior performance in place recognition, loop closure detection, and global localization.
@@ II Related Work
This section briefly overviews LiDAR-based global localization methods in the literature. Following existing surveys [27, 28], we categorize related works into four groups based on the integration degree of place recognition and pose estimation: 1) place recognition-only, that retrieves the most similar place using global descriptors; 2) global pose estimation, that directly estimates the global poses without retrieving places; and 3) place recognition followed by local pose estimation, that first achieves place recognition and then estimates the robot pose via a pose estimator. 4) As using BEV image representation for LiDAR localization, we additionally review the image processing techniques for BEV images.
@@
1) Place recognition-only. These methods generate global descriptors for LiDAR scans and perform place recognition through descriptor retrieval. The pose of the retrieved place is then considered as the estimated pose. The key to place recognition is creating global descriptors that ensure similar scans are close in descriptor space while dissimilar ones are far apart. Early methods usually exploit the point statistics to represent the point cloud appearance. For example, M2DP [29] projects a point cloud to multiple 2D planes and generates a density signature for points in each plane. The singular value decomposition (SVD) components of the signature are then used to compute a global descriptor. Scan Context [30, 31] partitions the ground space into bins according to both azimuthal and radial directions and encodes the 2.5-D information within an image.
@@
Recently, learning-based place recognition has been a hot topic, and many methods have been proposed. PointNetVLAD [14] leverages a network to project each point into a higher dimension feature and then uses NetVLAD [23] to generate global features. To take advantage of more contextual information, PCAN [32] introduces the point contextual attention network that learns attention to task-relevant features. Both PointNetVLAD and PCAN cannot capture local geometric structures due to the independent treatment for each point. Thus, the following methods focus on extracting more distinctive local features considering the neighborhood information. LPD-Net [33] adopts an adaptive local feature module to extract the handcrafted features and uses a graph-based neighborhood aggregation module to discover the spatial distribution of local features. EPC-Net [34] improves LPD-Net by using a proxy point convolutional neural network. SOE-Net [35] introduces a point orientation encoding (PointOE) module. Minkloc3D [36, 37] uses sparse 3D convolutions in local areas and achieves state-of-the-art performance on the benchmark dataset. Recently, some works including SVT-Net [38], TransLoc3D [39], NDT-Transformer [40], and PPT-Net [41] leverage the transformer-based attention mechanism [42] to boost place recognition performance. Some methods explore the potential of the range image representation of point clouds. OverlapNet [43] uses the overlap of range images to determine whether two point clouds are at the same place and uses a Siamese network to estimate the overlap. OverlapTransformer [20] further uses a transformer architecture to learn rotation-invariant global features. CVTNet [22] combines range images and BEV images to perform matching. It transforms BEV images into a format similar to range images to achieve rotation-invariance. The range image-based methods usually have better generalization ability but suffer from the scale distortions of translation movements of point clouds.
@@
2) Global pose estimation. These methods utilize the local characteristics of point clouds through geometric measures to generate local descriptors, which are then directly matched with a global map to determine the global poses. For example, the fast point feature histograms (FPFH) [44] uses the relationship between the neighbors of interest points and encodes them into a histogram. The Signature of Histograms of OrienTations (SHOT) [45] builds a local reference frame and leverages the orientation distribution of the normals in local regions. Although these methods can align local LiDAR scans, they are sensitive to point cloud density and noise. To leverage the local descriptors more efficiently, the keypoints voting method [46] performs localization using the 3D Gestalt descriptors. This method relies on robust keypoint extraction, while repeatable 3D keypoint detection is still an open problem in the literature. Some studies [47, 48] directly extract conventional image descriptors such as ORB [12] and SIFT [13] from LiDAR images for place recognition. SegMatch [49] performs segmentation on LiDAR scans and builds a segment map. During online global localization, it extracts segments from the query LiDAR scan and matches them with the map. The following SegMap [50] further assigns the segment descriptors learned by a neural network to improve the matching performance. The main drawback of these methods is their reliance on traveling relatively long distances to extract distinctive segments by gathering multiple scans.
@@
3) Place recognition followed by local pose estimation. These approaches have been commonly used in visual global localization. However, it is not widely adopted in LiDAR global localization because few LiDAR local features have reached the same level of maturity as visual features like ORB and SIFT. To tackle the problem, BVMatch [51] projects point clouds into BEV images and extracts handcrafted BVFT features from the images. It then uses the bag-of-words model [7, 8, 11] to generate global features. BoW3D [52] extracts local features directly from sparse point clouds. It also adopts the bag-of-words model to generate global descriptors and is expected to show moderate generalization ability. Another line of work adopts deep learning techniques. For example, DH3D [15] uses the 3D local feature encoder and detector to extract local descriptors. It embeds the descriptors to a global feature for place recognition and aligns the matched LiDAR pairs using RANSAC [25]. The following methods such as LCDNet [17], LoGG3D-Net [24], and LCRNet [18] use a similar strategy and unify place recognition and pose estimation into one framework. LCDNet extracts distinctive local features with pointvoxel-RCNN [53] and performs data association with the Sinkhorn algorithm [54]. LoGG3D-Net [24] introduces a local consistency loss to guide the network toward learning local features consistently across revisits. LCRNet [18] exploits novel feature extraction and pose-aware attention mechanism to precisely estimate similarities and 6-DoF poses between pairs of LiDAR scans. By jointly learning local and global descriptors, these methods show satisfactory global localization performance. However, they may not adapt well when the point clouds are out of the distribution of the training data. BTC [55] extracts the keypoints of a point cloud by projecting the points to planes. It generates triangle descriptors and develops an efficient matching strategy.
4) BEV-based methods. The primary challenge in processing BEV images is designing rotation-robust and discriminative features. Although some handcrafted features such as BVFT [51], HOPN [56], and RING [57] have shown promising performance, learning BEV features with neural networks remains rarely studied. Some BEV-based end-to-end networks [58, 59] work well for segmentation, but they do not extract features for localization. In this work, we propose the rotation equivariant module (REM) to address this gap. Notably, rotation equivariant networks have been used in other vision tasks such as object detection [60, 61] and image matching [62]. Compared to the rotation equivariance design in these methods and previous BEVPlace [26], our REM is explicitly designed to preserve the distinctiveness of deep features by directly leveraging CNN backbones within the architecture. This enables REM to capture richer and more robust features in a rotationally invariant manner.